Two angles are compliments and one lacks 8 degrees of being twice the other. What are the angles?
Thank you

Found 2 solutions by ewatrrr, london maths tutor:
Answer by ewatrrr(24785)

Hi
*Note: The sum of complementary degrees is 90 degress
Let c represent one of the complementary anlges.
According to the Question:(2c-8) will represent the other.
.
c + (2c-8) = 90
.
combine like terms and solve for x
3c - 8 = 90
.
c = 32 2/3 degrees
It's complement 57 1/3
.

Answer by london maths tutor(243)

Let one angle be x and the other angle be y
x + y = 180 degree (Two angles are compliments)
2y = x - 8 (one lacks 8 degrees of being twice the other)
x = 2y + 8
combining the above equations:
x + y = 180
(2y + 8) + y = 180
2y + y = 180 - 8
3y = 172
y = 57.3
Therefore, x = 180-57.33 = 122.67
Answer: angles are 122.67 and 57.33
Hope this is helpful.
For more examples, visit http://www.london-maths-tutor.co.uk/